Grade 12IIT JEE Entrance Exam

The number of waves made by a Bohr electron in one complete revolution in its 3rd orbit - (A) 1 (B) 2 (C) 3 (D) 4

Subham Panda
9 Years agoGrade 12

1 Answer

Sudhanshu Sahil
8 Years ago
MVR=nh/2πW(wavelength)=h/p (de Broglie)W=h/mvTherefore,(h/W) × R = 3h/2π (n=3 for 3rd orbit)3W = 2πR Which means no. Of waves is 3..